# Cost-Efficient Subjective Task Annotation and Modeling through Few-Shot Annotator Adaptation

### v1

@@ 1 Introduction
We introduce a unique dataset that enables the study of detecting moral content, an understudied subjective task, at a scale that was not possible before [^1] The dataset will be released as part of the accepted paper. The Moral Foundations Subjective Corpus (MFSC) is a collection of 2000 Reddit posts, each annotated by 24 annotators for moral content along with annotators’ responses to a range of psychological questionnaires (§4.1).
We use MFSC in conjunction with the Brexit Hate Dataset (Akhtar et al., 2021) to extensively study each component of our proposed framework. First, we empirically investigate the effect of pooling data from varying numbers of annotators in a multitask model (§5.1) and demonstrate that increasing the number of annotators does not improve annotator-level modeling in a multitask model. Second, we showcase the efficacy of our framework in capturing diverse annotator perspectives under budget constraints (§5.2). Third, in §5.3, we study the impact of various sample selection strategies. Our framework achieves a 4.3% increase in $F_{1}$ score with access to just 25% of the annotation budget in moral sentiment prediction. Furthermore, our results demonstrate a 6.7% improvement over the previous state-of-the-art in hate speech detection when using only 50% of the original annotation budget. Next, we show that our proposed framework consistently yields a more equitable model by minimizing the performance disparity across various annotators (§5.4). Specifically, on the lowest budget scenarios, our approach reduces the standard deviation in performance across annotators by 7.5% and 1.1% on hate speech detection and moral foundation classification, respectively. Finally, we extend our analysis to investigate whether the selection of the initial set of annotators in the first stage of our framework is related to the model’s performance (§5.5).
@@ 2 Related Work
Capturing the Perspectives: To capture annotator-level labels, Akhtar et al. (2020) proposed dividing annotators into groups based on similar personal characteristics and creating different sets of gold standards for each group. Kanclerz et al. (2022) and Deng et al. (2023) incorporated knowledge about annotators into their models to make them personalized. Davani et al. (2022) propose a multitask approach, modeling each annotators’ perspective as a subtask, while having a shared encoder across the subtasks. Baumler et al. (2023) and Wang and Plank (2023) propose active learning methods for reducing the budget of data collection by proposing methods for collecting samples based on model confidence and annotators’ disagreement. Casola et al. (2023) also proposes ensembling perspective-aware models based on their confidence.
@@ 3 Method
Problem Formulation: To formalize the task, suppose we have a set of annotators $\mathcal{A}=\{a_{1},...,a_{n}\}$ and input texts $X=\{x_{1},x_{2},...,x_{m}\}$ and their corresponding annotations $Y=\{y_{1},y_{2},...,y_{m}\}$. Let $D=\{D_{a_{i}}|a_{i}\in\mathcal{A}\}$ be the entire annotations and $D_{a_{i}}=\{X_{a_{i}},Y_{a_{i}}\}$ denote data collected from annotator $a_{i}$. Then the budget $B=|D|$ is defined as the total number of annotations collected. Let =$\mathcal{F}=\{f_{a_{i}}|a_{i}\in\mathcal{A}\}$ and $f_{a_{i}}$ denote the model capturing labels assigned by annotator $a_{i}$.
@@
Proposed Framework: We design our framework with two objectives: first, maximizing the average performance over all annotators. Second, minimizing the budget ($B$) required to achieve the first goal. The second objective allows us to increase the number of annotators’ perspectives ($|\mathcal{A}|$ ) captured with a given budget. Our framework design is based on two key intuitions. Firstly, as we show in Section 5.1, multitask learning, which has often been treated as the upper bound by previous work, does not improve in performance as the number of annotators grows. Secondly, even in subjective tasks, there exists a substantial number of texts on which annotators mostly agree, particularly when these texts are randomly drawn from a source. Therefore, obtaining many annotations on such instances is not beneficial in learning a new perspective. In line with these intuitions, our framework consists of two stages (Figure 1). In the first stage, we learn the commonalities between annotators through a multitask model $\mathcal{F}_{mtl}$. A crucial difference of our approach in comparison to previous multitask methods is that we only collect annotations from a small subset of annotators $\mathcal{A}_{mtl}\subset\mathcal{A}$. In the second stage, we learn the perspectives of new annotators $\mathcal{A}_{fs}=\mathcal{A}-\mathcal{A}_{mtl}$ with only a few shots. Specifically, we collect annotations for $k$ input texts $\mathcal{S}(X)\subset X$, where $\mathcal{S}$ is a sampling function that ideally helps in capturing patterns specific to individual annotators’ perspectives. Let $D^{fs}_{a_{i}}=\{(x,y_{ai})|x\in\mathcal{S}(X)\}$ and $|D^{fs}_{a_{i}}|=k<<|D_{a_{i}}|$. We initialize $\mathcal{F}_{\mathcal{A}_{fs}}$ with $\mathcal{F}_{mtl}$ and train it on $D^{fs}_{a_{i}}$.
@@
Sampling Function ($\mathcal{S}$): We explore four different sampling functions: 1) $\mathcal{S}_{rand}$: selects a random sample for each annotator 2) $\mathcal{S}_{mv}$: selects a balanced sample determined by the majority vote of the annotators. For a set of annotators $\mathcal{A}_{mtl}$, we calculate the majority vote among these annotators and select $k$ samples that have an equal number of each label based on that majority vote. 3) $\mathcal{S}_{dis}$ selects the samples from $\mathcal{A}_{mtl}$ with highest disagreement score, and 4) $\mathcal{S}_{bal}$ acts as an oracle, selecting a balanced sample based on a specific annotator’s label, not the majority vote. Therefore, if we have a new annotator, $\mathcal{S}_{bal}$ would select a balanced sample based on the annotations of that specific annotator. One frequent challenge in some subjective tasks is the heavy imbalance in class frequencies. Hence, we chose $\mathcal{S}_{mv}$ and $\mathcal{S}_{bal}$ to provide a more balanced sample to the few-shot model for each annotator. We added $\mathcal{S}_{dis}$ with the goal of providing samples that differentiate the individual annotator perspectives to the model. We use the “item disagreement” and “annotator disagreement” measure from Davani et al. (2023) to select samples in $\mathcal{S}_{dis}$.
@@ 4.1 Datasets
For a dataset to be well-suited for our experiments it should meet the following four constraints: 1) the dataset should pertain to a subjective topic, one that allows for different perspectives, 2) individual annotators’ labels should be available, 3) there should be enough annotators to allow experimentation with varying budgets, and 4) each sample should be annotated by all annotators. This will allow us to ensure any observed performance gains are due to the used methods and are not driven by the specific samples annotated by each annotator. Below, we describe the two datasets: the MFSC, which we collected specifically to enable research on subjective tasks, and the Brexit Hate Dataset.
Moral Foundations Subjective Corpus (MFSC): We introduce a new dataset, Moral Foundations Subjectivity Corpus (MFSC), consisting of 2000 Reddit posts annotated by 24 annotators for moral sentiment based on the Moral Foundations Theory (MFT; Graham et al., 2013; Atari et al., 2022). Morality is widely acknowledged to be a subjective concept, strongly influenced by cultural backgrounds Graham et al. (2016), one that has not been explored much in the NLP community. We asked annotators to label each text for the specific moral concern (i.e., Purity, Harm, Loyalty, Authority, Proportionality or Equality), and if one of the concerns is chosen, we set the label as moral, else as non-moral; This dataset was collected following the same procedure as Trager et al. (2022). This dataset also contains additional metadata information, such as confidence for each instance using a 3-level measure (i.e., confident, somewhat confident, and not confident). We also collected annotator responses for the “Big Five Inventory-2-Short” questionnaire (Soto and John, 2017). MFSC provides an opportunity to explore the subjective nature of morality. The substantial number of annotators in this dataset along with questionnaire responses enables future researchers to investigate the modeling of subjective tasks on a larger scale. The demographics of the annotators is shown in Appendix A.1. In our experiments, we use the annotations for the moral label. In addition, we evaluate our framework on Care label which is discussed in Appendix B.2.
@@
Brexit Hate dataset: Hate speech detection has become one of the primary subjective tasks studied in the NLP community (Akhtar et al., 2019; Sang and Stanton, 2022; Sap et al., 2021). The Brexit Hate dataset (Brexit) proposed by Akhtar et al. (2021), consists of 1,120 English tweets collected with keywords related to immigration and Brexit. The dataset was annotated with hate speech (in particular xenophobia and islamophobia), aggressiveness, offensiveness, and stereotype, by six annotators belonging to two distinct groups: a target group of three Muslim immigrants in the UK, and a control group who were researchers with Western background. For our experiments, we use the overall hate label. Additional dataset statistics can be found in Appendix A.
@@ 4.2 Experiment Setup
$=\frac{\sum_{a_{i}\in\mathcal{A}_{\text{mtl}}}|D_{a_{i}}|}{\sum_{a_{i}\in\mathcal{A}}|D_{a_{i}}|}\times B$
$=\frac{|\mathcal{A}_{mtl}|}{|\mathcal{A}|}\times B$
@@ 4.3 Implementation Details
We employ RoBERTa-base as the base model for our methods (Liu et al., 2019). All multitask models undergo hyperparameter tuning, specifically for learning rate and weight decay, and are trained for 5 epochs. We employ a weighted random sampler for the Brexit dataset to account for the imbalance in the labels of each annotator. Our training batch size is 128, and we use AdamW optimizer. The best hyperparameters of the best-performing model are then applied in the subsequent fine-tuning step. Few-shot models are trained for 50 epochs, during which all model parameters are updated. All models’ training is repeated with three random seeds. For the Brexit dataset, we adopt the predefined train, validation, and test splits provided within the dataset  [^2] https://le-wi-di.github.io/. For the MFSC dataset, we allocate 80% for training, 10% for validation, and the remaining 10% for testing.
@@ 5 Results and Analysis
5.1 Few Annotators are Enough
$metric=F_{1}^{overall}\uparrow$ Brexit MFSC

### v2

@@ 1 Introduction
We introduce a unique dataset that enables the study of detecting moral content, an understudied subjective task, at a scale that was not possible before [^1] The dataset will be released as part of the accepted paper. The Moral Foundations Subjective Corpus (MFSC) is a collection of 2000 Reddit posts, each annotated by 24 annotators for moral content along with annotators’ responses to a range of psychological questionnaires (§4.1).We use the MFSC dataset in conjunction with the Brexit Hate Dataset (Akhtar et al., 2021) to extensively study each component of our proposed framework. We evaluate our framework on three models: RoBERTa-Base, RoBERTa-Large, and Llama-3. In section 5.1, we demonstrate the efficacy of our framework in capturing diverse annotator perspectives under budget constraints. Our framework achieves a 4% increase in $F_{1}$ score with access to just 50% of the annotation budget in hate speech detection, and a 2% gain in moral sentiment detection with as little as 25% of the original annotation budget. Furthermore, we evaluate the efficiency of our framework in scaling to more annotators, i.e., incorporating a new annotator into an already existing annotated dataset and model through our second-stage few-shot adaptation. Our results show an $F_{1}$ score gain of 9% and 4% in the Brexit and MFSC datasets, respectively, demonstrating the scalability of our framework. Next, in section 5.2, we show that our proposed framework yields a more equitable model by minimizing performance disparity across annotators. Specifically, in the lowest budget scenarios, our approach reduces the standard deviation of the performance across annotators by 7% in hate speech detection and by 1% in moral sentiment classification. Finally, in section 5.3, we extend our analysis to investigate whether the selection of the initial set of annotators in the first stage of our framework affects the model’s performance.
@@ 2 Related Work
Capturing the Perspectives: One method for learning directly from crowd annotations is using soft loss, where the probability distributions of item labels are used as soft targets in a loss function (Peterson et al., 2019). However, this approach does not provide individual predictions for annotators, making it unsuitable for subjective tasks that require such specificity. To capture annotator-level labels, Akhtar et al. (2020) proposed dividing annotators into groups based on similar personal characteristics and creating different sets of gold standards for each group. Kanclerz et al. (2022) and Deng et al. (2023) incorporated knowledge about annotators into their models to make them personalized. Davani et al. (2022) propose a multitask approach, modeling each annotators’ perspective as a subtask, while having a shared encoder across the subtasks. Baumler et al. (2023) and Wang and Plank (2023) propose active learning methods for reducing the budget of data collection by proposing methods for collecting samples based on model confidence and annotators’ disagreement. Casola et al. (2023) also proposes ensembling perspective-aware models based on their confidence.
@@ 3 Method
Problem Formulation: To formalize the task, suppose we have a set of annotators $\mathcal{A}=\{a_{1},...,a_{n}\}$ and input texts $X=\{x_{1},x_{2},...,x_{m}\}$ and their corresponding annotations $Y=\{y_{1},y_{2},...,y_{m}\}$. Let $D=\{D_{a_{i}}|a_{i}\in\mathcal{A}\}$ be the entire annotations and $D_{a_{i}}=\{X_{a_{i}},Y_{a_{i}}\}$ denote data collected from annotator $a_{i}$. Then the budget $B=|D|$ is defined as the total number of annotations collected. Let =$\mathcal{F}=\{f_{a_{i}})|a_{i}\in\mathcal{A}\}$ and $f_{a_{i}}$ denote the model capturing labels assigned by annotator $a_{i}$.
@@
Proposed Framework: We design our framework with two objectives: first, maximizing the average performance over all annotators. Second, minimizing the budget ($B$) required to achieve the first goal. The second objective allows us to increase the number of annotators’ perspectives ($|\mathcal{A}|$ ) captured with a given budget. Our framework design is based on two key intuitions. Firstly, as shown in Figure 3, multitask learning (the orange line), which has often been treated as the upper bound by previous work, does not always improve in performance as the number of annotators grows. Secondly, even in subjective tasks, there exists a substantial number of texts on which annotators mostly agree, particularly when these texts are randomly drawn from a source. Therefore, obtaining many annotations on such instances is not beneficial in learning a new perspective. In line with these intuitions, our framework consists of two stages (Figure 1). In the first stage, we learn the commonalities between annotators through a multitask model $\mathcal{F}_{mtl}$. A crucial difference of our approach in comparison to previous multitask methods is that we only collect annotations from a small subset of annotators $\mathcal{A}_{mtl}\subset\mathcal{A}$. In the second stage, we learn the perspectives of new annotators $\mathcal{A}_{fs}=\mathcal{A}-\mathcal{A}_{mtl}$ with only a few shots. Specifically, we collect annotations for $k$ input texts $\mathcal{S}(X)\subset X$, where $\mathcal{S}$ is a sampling function that ideally helps in capturing patterns specific to individual annotators’ perspectives. Let $D^{fs}_{a_{i}}=\{(x,y_{ai})|x\in\mathcal{S}(X)\}$ and $|D^{fs}_{a_{i}}|=k<<|D_{a_{i}}|$. We initialize $\mathcal{F}_{\mathcal{A}_{fs}}$ with $\mathcal{F}_{mtl}$ and train it on $D^{fs}_{a_{i}}$.
@@
Sampling Function ($\mathcal{S}$): We explore four different sampling functions: 1) $\mathcal{S}_{rand}$: selects a random sample for each annotator 2) $\mathcal{S}_{mv}$: selects a balanced sample determined by the majority vote of the annotators. For a set of annotators $\mathcal{A}_{mtl}$, we calculate the majority vote among these annotators and select $k$ samples that have an equal number of each label based on that majority vote. 3) $\mathcal{S}_{dis}$ selects the samples from $\mathcal{A}_{mtl}$ with highest disagreement score, and 4) $\mathcal{S}_{bal}$ acts as an oracle, selecting a balanced sample based on a specific annotator’s label, not the majority vote. Therefore, if we have a new annotator, $\mathcal{S}_{bal}$ would select a balanced sample based on the annotations of that specific annotator. One frequent challenge in some subjective tasks is the heavy imbalance in class frequencies. Hence, we chose $\mathcal{S}_{mv}$ and $\mathcal{S}_{bal}$ to provide a more balanced sample to the few-shot model for each annotator. We added $\mathcal{S}_{dis}$ with the goal of providing samples that differentiate the individual annotator perspectives to the model. We use the “item disagreement” and “annotator disagreement” measures from Davani et al. (2023) to select samples in $\mathcal{S}_{dis}$.
@@ 4.1 Datasets
We run experiments on two datasets annotated for subjective tasks: Brexit Hate dataset (Akhtar et al., 2021) and the Moral Foundations Subjective Corpus (MFSC), which we created as part of this work to explore this less-studied subjective task. Both datasets contain per-annotator labels for instances, with every instance being annotated by all annotators. This ensures that any observed performance gains are attributed to our method, rather than the specific samples annotated by each annotator. Additionally, we evaluate our framework on the Gab Hate Corpus (GHC; Kennedy et al., 2018), where the number of annotations by different annotators varies. Detailed experiments and results for this dataset are presented in Appendix C.2.
Moral Foundations Subjective Corpus (MFSC): We introduce the Moral Foundations Subjective Corpus (MFSC), a new dataset consisting of 2000 Reddit posts annotated by 24 annotators for moral sentiment based on the Moral Foundations Theory (MFT; Graham et al., 2013; Atari et al., 2023). Morality, being a subjective concept heavily influenced by cultural backgrounds Graham et al. (2016), has not been extensively explored in the NLP community.
Each sample in the MFSC is annotated with a binary label indicating moral sentiment: 1 if the sentence pertains to morality and 0 if it does not. We utilize this binary moral/non-moral label in our experiments. Additionally, we have collected more fine-grained labels of morality, which are detailed in the Appendix A. Examples of the dataset and their annotations for moral sentiment are presented in Table 1. The demographics of the annotators are provided in Appendix A.1.
MFSC examples a1 a2 a3 a4 a5 a6
You’re a horrible person, and deserve the same thing to happen to you. 1 1 1 1 1 1
As an expat Brit, I was moved: What a brilliant unifying speech. Here’s fingers crossed for you USA. 1 0 0 1 1 1
That meal is insane compared to what we got. Don’t think we ever had fresh veg/fruit. 0 0 0 0 0 0
Table 1: Examples from the MFSC dataset with binary labels for moral sentiment. The examples show the labels provided by 6 out of 24 annotators.
@@
Brexit Hate dataset: Hate speech detection has become one of the primary subjective tasks studied in the NLP community (Akhtar et al., 2019; Sang and Stanton, 2022; Sap et al., 2021). The Brexit Hate dataset (Brexit) introduced by Akhtar et al. (2021), consists of 1,120 English tweets collected with keywords related to immigration and Brexit. The dataset was annotated with hate speech (in particular xenophobia and islamophobia), aggressiveness, offensiveness, and stereotype, by six annotators belonging to two distinct groups: a target group of three Muslim immigrants in the UK, and a control group who were researchers with Western background. For our experiments, we use the overall hate label.
Table 2 provides the datasets’ statistics, including Fleiss’s kappa (Fleiss, 1971), which measures the inter-annotator agreement. The low agreement values highlight the subjective nature of these tasks. Furthermore, the ’%Pos.’ column in Table 2 shows the class imbalance in the Brexit dataset and the scarcity of positive class annotations. For example, in the Brexit dataset, only 12% of samples, on average, were labeled as "Hate".
[…]
Brexit $1120$ $6$ $0.34$ $12.86$
MFSC (Moral) $2000$ $24$ $0.26$ $63.69$
[…]
@@ 4.2 Experiment Setup
$=\frac{\sum_{a_{i}\in\mathcal{A}_{\text{mtl}}}|D_{a_{i}}|}{\sum_{a_{i}\in\mathcal{A}}|D_{a_{i}}|}\times B=\frac{|\mathcal{A}_{mtl}|}{|\mathcal{A}|}\times B$
@@ 4.3 Implementation Details
We evaluate our framework using three base models: RoBERTa (both base and large versions) (Liu et al., 2019), and Llama-3 (Touvron et al., 2023).
Roberta Models: All multitask models undergo hyperparameter tuning for learning rate and weight decay (see Appendix D.1) and are trained for 5 epochs. The best model is selected based on the validation $F_{1}$ score, and its optimal hyperparameters are also applied in the few-shot stage. All models converge within 5 epochs for MTL and 50 epochs for few-shot learning. To ensure robustness, experiments are repeated with three random seeds.
Llama-3: We use Llama-3-8B and employ LoRA (Low-Rank Adaptation; Hu et al., 2021) for fine-tuning. We conduct hyperparameter tuning for LoRA parameters, in addition to learning rate and weight decay. In the second stage of our framework, we use the same set of hyperparameters determined in the first stage for few-shot adaptation. The hyperparameters used in our MTL training, along with other variables, are shown in Table 7.
For all models, we use the AdamW optimizer. For the Brexit dataset, we utilize predefined train, validation, and test splits provided within the dataset [^2] https://le-wi-di.github.io/, and we employ a weighted random sampler to account for the imbalance in the labels of each annotator. For the MFSC dataset, we allocate 80% for training, 10% for validation, and the remaining 10% for testing.
In the few-shot stage, we experiment with four different values of $k$ (16, 32, 64, and 128). We report the results for $k=128$ in the next section, while the experiments for other values of $k$ are provided in Appendix D.
@@ 5 Results and Analysis
5.1 Towards Better Performance with Less Annotation Budget
Figure 2: Overall $F_{1}$ score ($F_{1}^{overall}$) of our framework compared to the baseline across all three base models on both datasets. We observe a 3.8% performance gain with only 50% of the annotation budget on Brexit dataset, and 2.24% gain with 25% of the annotation budget on MFSC dataset, on the best performing base models.
Figure 2 shows the overall $F_{1}$ scores of our framework for two datasets across varying budgets, evaluated using three different base models. We observe that our framework consistently outperforms the baseline, particularly at lower budget levels. More importantly, our method surpasses the baseline trained with 100% of the budget using as little as 25% of the original budget across all three base models, demonstrating its model-agnostic efficacy.
Specifically, at the lowest budget level in the Brexit dataset, our framework with balanced sampling ($\mathcal{S}_{bal}$) achieves performance gains of 5%, 14%, and 5% compared to the baseline when trained with RoBERTa-Base, RoBERTa-Large, and Llama-3, respectively. Compared to the full budget, our method shows a gain of 3.8% with RoBERTa-Large and 3.38% with Llama-3 using only 50% of the original budget, and a gain of 4.34% with RoBERTa-Base using 66% of the budget.
For the MFSC dataset, our framework, regardless of the sampling method, outperforms the baseline across all budget levels with the RoBERTa-Base and RoBERTa-Large models. Additionally, with the Llama-3 model at 25% of the budget, our method has 4% gain compared to baseline.
These findings demonstrate the success of our framework in achieving its dual objectives: enhancing performance across all annotators while reducing annotation budget requirements. We also conduct an ablation study by omitting the first MTL stage and employing random few-shot sampling for each annotator. Additionally, we compare our framework with more baselines (see Appendix B).
Incorporating a New Annotator: The second stage of our framework suggests that few-shot adaptation not only allows us to integrate a new annotator into an already existing model with minimal budget, but also maintains the annotator’s performance. To validate this ability, in Figure 3 we compare the performance of the second stage of our framework ($F_{1}^{fs}$) with the baseline.
For the Brexit dataset, $F_{1}^{fs}$ scores exceed the baseline across all base models by 9.28%, 8.47%, and 5.07%, respectively. The balanced sampling ($\mathcal{S}_{bal}$) method consistently performs well across all models. Similarly, for the MFSC dataset, our framework achieves higher $F_{1}^{fs}$ scores regardless of sampling method, except with Llama-3 model.
Overall, our results on both datasets show that the few-shot stage of our framework results in models that outperform the multitask learning baseline.
Figure 3: Few-shot $F_{1}$ score ($F_{1}^{fs}$) of our framework compared to the baseline across all three base models on both datasets. We observe a 8.47% performance gain with 83% of the annotation budget on Brexit dataset, and 4.37% gain with 25% of the annotation budget on MFSC dataset, on the best performing base models.
Base Model Comparison: Generally, the RoBERTa models perform better than the Llama-3 model in MTL setting. Llama-3 model, despite undergoing the most hyperparameter search and utilizing the most GPU hours to find optimal parameters, still performs significantly poorer than the other two models, especially when fine-tuned in a few-shot setting. A potential reason for this disparity is that larger models like Llama-3, while generally more capable, require extensive hyperparameter tuning to optimize their performance. Additionally, they have stronger biases, making it more challenging to adapt them to different perspectives (Naveed et al., 2023; Liu et al., 2023).
5.2 Reduced Performance Disparities across Annotators
Ensuring a comprehensive representation of annotators’ viewpoints is crucial in modeling subjective tasks. To achieve this goal, a critical criterion is to create models that not only improve the aggregated performance but also demonstrate fair and equitable performance across all annotators. For example, if the $F_{1}$ scores of one model for two annotators are 0.6 and 0.8, respectively, while the second model scores 0.7 for both annotators, the latter is considered a better model. Although the average performance is the same for both models, the first model has a disparate negative impact on the first annotator. This is important because performance disparities among social groups (in our case annotators) can lead to biased models, limiting the system’s ability to accurately reflect diverse perspectives and potentially perpetuating inequalities in the outputs of subjective tasks (Buolamwini and Gebru, 2018). Merely relying on aggregated performance measures, such as the average across all annotators, fails to provide a comprehensive understanding of how well the model captures the varying perspectives of different annotators. For instance, it remains unclear whether the average performance improves because the approach better captures the perspectives of all the annotators or only a subset of them. Hence, we look into the standard deviation of performance across all annotators as a measure of performance disparity: $d=\sqrt{\frac{1}{N-1}\sum_{i=1}^{N}(F^{a_{i}}_{1}-\overline{F^{overall}_{1}})^{2}}$. Lower standard deviations indicate more equitable models.
As shown in Table 3, our approach results in lower performance disparities ($d$) compared to the MTL baseline regardless of the base model, across all budgets for the MFSC dataset. For the Brexit dataset, this improvement is observed at lower budgets (50% and 66%). Among the various sampling strategies, the balanced sampling strategy ($S_{bal}$) consistently results in lower $d$ for MFSC dataset. When comparing different base models, the lowest $d$ is achieved using RoBERTa-Base model. Specifically, for the MFSC dataset, there is a 1.1% reduction in $d$ at 25% of the budget compared to the baseline, and for the Brexit dataset, there is a 7.5% reduction in $d$ at 50% of the budget. Figure 4 visualizes this model’s performance in comparison to the MTL baseline for each annotator. Notably, our framework improves performance for annotators in the non-Western control group (i.e., the first three annotators) while maintaining the performance of the remaining annotators.
Overall, these findings suggest that our proposed framework not only improves the overall performance of all annotators but also yields models that are more fair and equitable.
     $d\downarrow$ Brexit MFSC